OpenStudy (anonymous):

If an area of an isosceles right triangle is 36, what is the perimeter?

OpenStudy (anonymous):

\[A) 12+12\sqrt{2} \] \[B) 12+6\sqrt{2} \] \[C) 6+12\sqrt{2}\] \[D) 18\sqrt{2}\] \[E) 12\sqrt{2}\]

OpenStudy (anonymous):

Any other relations u can provide??

OpenStudy (anonymous):

ummm nope

OpenStudy (anonymous):

i guess B

OpenStudy (anonymous):

M not able to define any relation....sry on dat!!

OpenStudy (anonymous):

well thanx anyway !!!! :)

OpenStudy (anonymous):

hmmm....k

OpenStudy (anonymous):

@genius12 can you help???

OpenStudy (anonymous):

@GoldEverything help??

OpenStudy (anonymous):

@Luigi0210 help??

OpenStudy (goldeverything):

how could i be of service :)

OpenStudy (anonymous):

hi!!Cant seem to find any relation in the question,can ye??

OpenStudy (anonymous):

Oh shoot!!its a r8 triangle

OpenStudy (anonymous):

I m xtremely sry guys,dint read it properly,was in hurry!!

OpenStudy (anonymous):

Sorry to @genius12 @Luigi0210 @GoldEverything !!!!

OpenStudy (anonymous):

im still thinking b tho....

OpenStudy (anonymous):

So here it is.....its an isosceles r8 triangle....therefore the two sides which are equal have to be both sides other than hypotenuse!!

OpenStudy (anonymous):

so area =(1/2).a.a=36,where a=adjacent or opposite side,get it??

OpenStudy (anonymous):

k

OpenStudy (anonymous):

from this find a,then find the hypotenuse,then the perimeter...cheers!!

OpenStudy (anonymous):

hypotenuse=(a^2+a^2)^(1/2)

OpenStudy (anonymous):

=a2^(1/2)

OpenStudy (anonymous):

Wil dat help @ineedhelpnow08 ??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

k....then...happy to help!!